OpenStudy (anonymous):

A card is drawn from a deck of 52. Find the probability of drawing a red card (hearts or diamonds) OR a face card (jack, queen, or king)

OpenStudy (anonymous):

How many red cards in a deck? How many face cards in a deck? How many are both?

OpenStudy (anonymous):

26 red cards, 12 face cards, and 6 are both

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

go ahead @tomdelacruz u r correct

OpenStudy (anonymous):

4/13. Got it.

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

\[\bf P(A∪B)=P(A)+P(B)-P(A∩B)\] @tomdelacruz From what you mentioned about "6 being both", we note that the two events are not mutually exclusive. So just add the probabilities of each and subtract the probability of getting a card that is both. That's it.

OpenStudy (uri):

There are 26 red cards and 12 face cards.Total cards in a deck=52 So 26/52 and 12/52 26/52=1/2 12/52=3/13 1/2+3/13=19/26.

OpenStudy (anonymous):

@uri however those events are not mutually exclusive so you must subtract their intersection (6/52) which will result in 32/52 = 4/13

OpenStudy (uri):

Ye,I was just trying.I don't even have this subject :p @Paynesdad

OpenStudy (anonymous):

@uri not bad w/o taking the course.

OpenStudy (anonymous):

@Paynesdad 32/52 is not = to 4/13

OpenStudy (anonymous):

32/52 = 8/13

OpenStudy (anonymous):

lol you are right ... my bad... must be getting tired.

OpenStudy (uri):

@Paynesdad Thanks.Can you tell me where i went wrong?

OpenStudy (anonymous):

Well you counted some cards twice. You counted the 26 red cards and the 12 face cards. However, 6 of the face cards are also red cards so the total number of either red cards or face cards is 26(red cards total) + 12(face cards total) - 6(cards that are both) = 32 out of 52 total. 32/52 as yahoo pointed out is 8/13 (not 4/13)

OpenStudy (uri):

Oh yeaaaaaaaaah! :D

OpenStudy (uri):

You're in which grade?

OpenStudy (anonymous):

LOL ... I wish. I am an engineer now. I taught hs math for 15 years before that.

OpenStudy (anonymous):

And coached wrestling...(sorry had to add that!... Save Olympic Wrestling)

OpenStudy (uri):

I'm only 16.Next time you're teaching me Probability.

OpenStudy (anonymous):

Anytime.